If the demand function is Q=110-20p and the supply function is Q 30+30p, what are the equilibrium price and quantity? The equilibrium price is S per unit. (Enter your response rounded to two decimal places.) The equilibrium quantity is units. (Enter your response rounded to one decimal place.)
Answer #1

Answer
the equilibrium is at Qd=Qs
110-20P=30+30P
50P=80
P=50/80=0.625
Q=20+30*(50/80)
=38.75
The equilibrium price is $0.63 and quantity is 38.8 units